OpenStudy (anonymous):

A square has an area of 16 cm². Calculate the length of its diagonal to the nearest 10th of a centimeter. Page 411 will be helpful after you find the length of the side of the square. a. 6.4 cm b. 5.7 cm c. 8 cm d. 4 centimeters e. 64 cm

OpenStudy (anonymous):

If the area is 16, each side is 4. Use the pathogen theorem 4^2+4^2=diagonal^2. Works out to about 5.7.

jhonyy9 (jhonyy9):

4sqrt2 =4*1,41 =5,64 than b.is right

OpenStudy (anonymous):

 You can calculate the diagonal with pythagoras: \[\sqrt{(4^2 + 4^2)} = \sqrt{32} \approx 5.7\]